# TTD: Text-Tag Self-Distillation Enhancing Image-Text Alignment in CLIP to Alleviate Single Tag Bias

### v1

@@ Abstract
We identify a critical bias in contemporary CLIP-based models, which we denote as single tag bias. This bias manifests as a disproportionate focus on a singular tag (word) while neglecting other pertinent tags, stemming from CLIP’s text embeddings that prioritize one specific tag in image-text relationships. When deconstructing text into individual tags, only one tag tends to have high relevancy with CLIP’s image embedding, leading to an imbalanced tag relevancy. This results in an uneven alignment among multiple tags present in the text. To tackle this challenge, we introduce a novel two-step fine-tuning approach. First, our method leverages the similarity between tags and their nearest pixels for scoring, enabling the extraction of image-relevant tags from the text. Second, we present a self-distillation strategy aimed at aligning the combined masks from extracted tags with the text-derived mask. This approach mitigates the single tag bias, thereby significantly improving the alignment of CLIP’s model without necessitating additional data or supervision. Our technique demonstrates model-agnostic improvements in multi-tag classification and segmentation tasks, surpassing competing methods that rely on external resources. Code is available at https://github.com/shjo-april/TTD.
@@
Figure 1: In existing image-text alignment research, single tag bias is evident where the image and text embeddings tend to concentrate solely on a single tag. (a) Single tag bias in image-text relationships: When examining the similarity map between the image and text, it is evident that only a single tag inside the red box is activated, disregarding other elements mentioned in the text. (b) Single tag bias in image-tag relationships: Even when observing the similarity values between each tag in the text and the image, a high value is assigned only to the single tag, while other tags that describe the image exhibit similar values to insignificant tags. We used TCL [4] for the analysis.
@@ 1 Introduction
Research on fully supervised classification/segmentation is widely applied in the field of computer vision. Various weakly/semi-supervised and few-shot methods have emerged to reduce labeling costs, but there has been active research in image-text alignment following CLIP [31]. Despite being trained solely on images and text (i.e., weak annotation), CLIP can be applied to downstream tasks that require stronger annotations, such as labels and masks. This study aims to improve CLIP performance using only weak information, which is in line with CLIP’s original design intent.
@@
As a major existing limitation to achieve this, we discover that CLIP-based models exhibit monofocus bias, where they tend to overemphasize a particular context of image-text relationship. As observed in Fig. 1(a), the similarity map between image and text reveals activation only in the region related to a single tag/word (i.e., chameleon), while neglecting others (i.e., branch and green). We termed this observed encoder output issue as single tag bias, indicating a flaw in CLIP’s pre-trained text encoder, which over-focuses on one tag, ignoring others.
@@
Addressing this issue involves creating accurate foreground regions for every relevant tag, not just the biased one, to fine-tune the encoder to recognize their collective areas. It requires identifying all relevant tags in the text. However, testing each tag through CLIP’s text encoder shows persistent single tag bias, highlighting the challenge of extracting genuine tags solely with image-text pairs, as seen in Fig. 1(b). This underlines the struggle of standard CLIP-based tag scoring due to the image encoder’s reduction of pixel data to a singular tag focus, blocking the recognition of other image-relevant tags directly from text, illustrated by the pixels’ overemphasis on one tag (i.e., jacuzzi in Fig. 2(a)), skewing the image’s global (pooling-based) representation towards it. Even when scoring based on the proportion of pixels similar to each tag within the image as in segmentation, it still fails to lower scores for irrelevant tags, thus being unable to correctly extract meaningful tags from the text (see Sec. 0.B.1).
@@
Figure 2: High-level overview of our method. (a) Tag selection by pixel-tag scoring: The pooled global image embedding predominantly reflects information about a single tag, which is “jacuzzi” in this example, due to the overactivation of a single tag in pixel embedding. (The colored similarity map in the figure shows the most relevant tag for each pixel.) By employing the maximum cosine similarity value between the tag and its most correlated pixel as a score, we minimize the impact of single tag bias observed in the image encoder. (b) Text-tag self-distillation: The similarity map between text and image demonstrates only a single tag, whereas it should represent all meaningful tags in the relationship between image and text. To alleviate the bias, we train the map between the image and text to align with the union of maps between the image and the pseudo tags obtained in (a), thus enhancing the image-text alignment.
@@
To address this, we shift from standard global pooling to a new method focusing on the pixel index nearest to each tag candidate. This is based on the finding that pixels closely correlated with a specific tag more accurately pinpoint image segments, as shown in Fig. 2(a). By using the closest pixel’s similarity to the tag as the measure for image-tag similarity, we counteract the bias, thus broadening the encoder’s tag representation capacity.
@@
We used pseudo tags from our scoring to remove unnecessary tag masks, creating valid foreground masks for all image-relevant tags. This led to a new self-distillation-based CLIP fine-tuning method, assuming optimal image-text relationships encompass all pseudo-tag contexts (Fig. 2(b)). We create a pseudo label, a union of the similarity maps between the image and all pseudo tags, making CLIP learn this as the real image-text map. This process enabled the model to recognize all related tag regions, not just one, improving CLIP’s image-text alignment without additional data or annotations.
@@
2 Related works
CLIP [31] stands out as the most prominent model for image-text alignment. It possesses a remarkable ability to comprehend various free-form texts, making it widely utilized in zero-shot/open-vocabulary scenarios. Consequently, various CLIP variants demonstrate comparable performance to models trained with task-specific annotations across many downstream tasks, such as classification and segmentation, even if they are trained with only image and text data.
Table 1: Comparison of our method with existing works
@@ 2.2.1 Utilization of External NLP models
Recent studies [23, 30, 38, 39, 24] have focused on extracting tags from text for learning, relying on external NLP models (Fig. 3(b)). While some studies utilize NLP parsers [26, 16, 1] to extract nouns, others leverage Large Language Models (LLM) like Vicuna [7]. However, these tag selection techniques entirely disregard image information, resulting in two issues as shown in Fig. 5: 1) Extracting image-irrelevant tags. 2) Overlooking image-relevant tags, particularly non-nouns. However, our method is free from such issues since we only utilize the image and text for tag selection.
@@ 2.2.2 Additional Filtering using Image Information
Some models [37, 9] additionally consider the image to score tags. However, they utilize globally pooled image representation, suffering from single tag bias (Fig. 1(b)). CoCu [37] additionally introduces scoring based on retrieved image relevancy, but this depends on the images retrieved, where tag-relevant images lead to a low score and vice versa, and also requires high computational cost (see Sec. 0.B.1). Also, both models rely on external NLP models and fail to address the omission of image-relevant tags.
@@
(a) Tag Selection by Pixel-Tag Scoring
(b) Method Overview
Figure 4: Overall framework of our method. (a) Tag selection by pixel-tag scoring: Utilizing the similarity with the most correlated pixel as scoring, we identify a significant gap between tags related to the image and irrelevant ones. This gap guides us in selecting tags with the largest score gap as the threshold. (b) Method overview: Our fine-tuning approach begins by obtaining pseudo tags through our tag selection process, which are then used to create an ideal heatmap representing the image-text relationship. We address single tag bias by distilling information between the ideal and actual heatmaps. Additionally, we integrate an auxiliary loss to learn the pixel-tag relationship. Throughout this process, both the image and text encoders are trained.
@@ 3.1 Tag Selection by Pixel-Tag Scoring
Given an image $I$ of size $H\times W$ and corresponding text $T$ composed of tags $\mathcal{T}=\{t_{1}\cdots t_{n}\}$, our objective is to extract tags, $\mathcal{T}_{gt}\subset\mathcal{T}$ that accurately represent the image-text relationship. (e.g. for the first image in Fig. 5, $\mathcal{T}_{gt}=\{t_{1},t_{2},t_{3},t_{6},t_{7}\}$.) Importantly, our approach relies solely on the image $I$ and the text $T$, without any external datasets or models.
Let $\mathcal{E}_{V},\mathcal{E}_{T}$ denote the image and text encoders, respectively, which provide pixel embeddings and global text embeddings: $\mathcal{E}_{V}(I)\in\mathbb{R}^{H\times W\times C},\mathcal{E}_{T}(T)\in\mathbb{R}^{C}$ where $C$ is the embedding dimension size of the encoder output and $H,W$ are height and width of the image. The global image embedding can be obtained by globally average pooling the pixel embedding: $\mathrm{gap}(\mathcal{E}_{V}(I))\in\mathbb{R}^{C}$.
@@
$s_{\mathrm{image}}(t_{i}):=\mathrm{sim}(\mathrm{gap}(\mathcal{E}_{V}(I)),\mathcal{E}_{T}(t_{i})),$ (1)
@@
Subsequently, we notice a substantial score gap between tags associated with the image and those not. As illustrated in Fig. 4(a), we arrange all tags according to $s_{\mathrm{pixel}}^{\mathrm{ours}}$ and compute the score difference between consecutive tags. We observed the significantly large score difference between meaningful tags and others at the boundary. Leveraging this observation, we extract pseudo tags with high scores using the largest score gap as the boundary: $\mathcal{T}_{pred}\subset\mathcal{T}$.
@@ 3.2 Text-Tag Self-Distillation
In this section, we introduce a self-distillation technique designed to mitigate single tag bias arising from text embedding (Fig. 1(a)). The pseudo tags $\mathcal{T}_{pred}$ obtained from our tag selection method (Sec. 3.1) encapsulate the relationship between the image and the text, enabling the transfer of knowledge from pseudo tags to the text. Specifically, we ensure alignment between the union of similarity maps of pseudo tags and the image with that of the text and image:
$\mathrm{sim}(x):=\mathrm{sim}(\mathcal{E}_{V}(I),\mathcal{E}_{T}(x))\in\mathbb{R}^{H\times W},$ (3)
$\mathcal{L}_{distill}=||\mathrm{sim}(T)-\bigcup_{t_{i}\in\mathcal{T}_{pred}}\mathrm{norm}(\mathrm{sim}(t_{i}))||^{2}_{2},$ (4)
where $\mathrm{sim}(x)$ is an abbreviation of the similarity map between the image $I$ and a text $x$, $\bigcup$ is the pixel-wise (element-wise) maximum operation, and $\mathrm{norm}$ is min-max normalization. Here, we apply a stop-gradient operation on the union similarity map to propagate the gradients only through the text-based one.
Additionally, we introduce an auxiliary loss function to reinforce the correlation between tags and the image. Tags that are not selected are deemed irrelevant to the image; thus, we drive their similarity map toward zero. Conversely, selected tags are trained to align with the discrete map to disregard unrelated image parts:
$\mathcal{L}_{tag}=\sum_{t_{i}\in\mathcal{T}}D(t_{i})\ \mathrm{,where}\ D(t_{i})=\begin{cases}||\mathrm{sim}(t_{i})-\mathrm{norm}(\mathrm{sim}(t_{i}))||^{2}_{2}&\mathrm{if}\ t_{i}\in\mathcal{T}_{pred}\\ ||\mathrm{sim}(t_{i})||^{2}_{2}&\mathrm{otherwise}\end{cases}.$ (5)
Similar to text-tag distillation loss, we apply a stop-gradient operation on the normalized masks.
@@
Some may question whether it is more effective to view segment-level relationships using external models like SAM [19] rather than examining pixel-level relationships. However, our method is unrestricted and versatile since we self-train the model utilizing only image and text information without additional supervision, such as ground-truth tag annotation or external models. We also provide supplementary experiments on advanced techniques utilizing external models in Sec. 0.B.5.
@@ 4 Experiments
4.0.1 Datasets and Implementation detail
In all experiments, we trained our model solely with image-text pairs without utilizing additional annotations like truth tag information for the text. We demonstrated its flexibility by evaluating its performance on two distinct public models [44, 4]. Furthermore, to cut computational costs, we utilized the LoRA [17] technique for all our training processes. We used the CC3M and CC12M datasets [34, 5] for training, applying an additional sample filtering strategy to reduce the use of largely uncorrelated image-text pairs.
To evaluate the performance of our tag selection method, we manually collected truth tag information for all image-text pairs of CC3M validation set [34]. Also, to see how much the single tag bias has eased, we generated the tag’s corresponding mask data for 680 image-text pairs. This was necessary because, to our knowledge, there is no existing dataset that provides a complete match of all truth tags and their corresponding masks with the text. We will make this information available through code release. Details are in Appendix 0.C.
Additionally, for assessing our method’s effectiveness in open-vocabulary semantic segmentation using tag inputs, we use six established benchmarks: PASCAL VOC (VOC) [12], PASCAL Context (Context) [29], COCO-Stuff (Stuff), COCO-Object (Object) [3], Cityscapes (City) [10], and ADE20k (ADE) [43]. Further evaluations on segmentation performance with referring datasets, alongside conventional open-vocabulary classification tasks, and specifics of these assessments are detailed in Sec. 0.A.1 and 0.B.
@@ 4.1 Main Results
(a) Comparison based on the use of NLP models
@@
(b) Comparison based on scoring methods

### v2

@@ Abstract
We identify a critical bias in contemporary CLIP-based models, which we denote as single tag bias. This bias manifests as a disproportionate focus on a singular tag (word) while neglecting other pertinent tags, stemming from CLIP’s text embeddings that prioritize one specific tag in image-text relationships. When deconstructing text into individual tags, only one tag tends to have high relevancy with CLIP’s image embedding, leading to biased tag relevancy. In this paper, we introduce a novel two-step fine-tuning approach, Text-Tag Self-Distillation (TTD), to address this challenge. TTD first extracts image-relevant tags from text based on their similarity to the nearest pixels then employs a self-distillation strategy to align combined masks with the text-derived mask. This approach ensures the unbiased image-text alignment of the CLIP-based models using only image-text pairs without necessitating additional supervision. Our technique demonstrates model-agnostic improvements in multi-tag classification and segmentation tasks, surpassing competing methods that rely on external resources. The code is available at https://github.com/shjo-april/TTD.
@@
Figure 1: Single Tag Bias. In existing image-text alignment research, single tag bias is evident where the image and text embeddings tend to concentrate solely on a single tag. (a) Single Tag Bias in Image-Text Relationships: When examining the similarity map between the image (i.e., pixels) and text, it is evident that only a region of a single tag inside the red box is activated, disregarding other tags mentioned in the text. (b) Single Tag Bias in Image-Tag Relationships: Even when observing the similarity values between each tag in the text and the image, a high value is assigned only to the single tag, while other tags that describe the image (e.g., “green” and “branch”) exhibit similar values to insignificant tags (e.g., “of” and “a”). We use TCL [4] for the analysis.
@@ 1 Introduction
Research on fully supervised classification/segmentation is widely applied in the field of computer vision. Various weakly/semi-supervised and few-shot methods have emerged to reduce labeling costs. Recently, there has been active self-supervised research in image-text alignment following CLIP [31]. Despite being trained solely on images and text (i.e., weak annotation), CLIP can be applied to downstream tasks that require stronger annotations, such as labels and masks.
@@
A major existing limitation we discovered with CLIP-based models is what we term as single tag bias, where they tend to overemphasize a particular context of image-text relationship. As observed in Fig. 1(a), the similarity map between image and text reveals activation only in the region related to a single tag/word (i.e., chameleon), while neglecting others (i.e., branch and green). This observed encoder output issue, single tag bias, indicates a flaw in CLIP’s pre-trained text encoder, which over-focuses on one tag, ignoring others.
@@
This paper aims to mitigate single tag bias in CLIP-based models using only image-text pairs, in line with CLIP’s original design intent. Addressing this issue involves creating accurate foreground regions for every relevant tag, not just the biased one, to fine-tune the model to recognize their collective areas. It requires identifying all relevant tags in the text. However, the similarity between the image and each tag embedding, obtained by inputting each tag through CLIP’s text encoder, shows persistent single tag bias as seen in Fig. 1(b). This underlines the struggle of standard CLIP-based tag scoring due to the image encoder’s reduction of pixel data to a singular tag focus, blocking the recognition of other image-relevant tags. (Fig. 2(a) shows the pixels’ overemphasis on one tag (i.e., “jacuzzi”), skewing the image’s global representation towards it.) Even when scoring based on the proportion of pixels similar to each tag within the image as in segmentation, it still fails to correctly extract meaningful tags from the text (see Sec. 0.D.1).
@@
Figure 2: High-level Overview of Our Method. (a) Tag Selection by Pixel-Tag Scoring: The global image embedding predominantly reflects information about a single tag, “jacuzzi” in this case, due to the overactivation of a single tag in pixel embeddings. The colored similarity map shows the most relevant tag for each pixel. (e.g., orange pixels have the highest similarity with “jacuzzi”.) We minimize the impact of single tag bias observed in the image embedding by employing cosine similarity between the tag and its most correlated pixel as a score. (b) Text-Tag Self-Distillation: The similarity map between image and text demonstrates only a single tag, whereas it should represent all relevant tags in the image-text relationship. To alleviate the bias, we train the image-text map to align with the union of maps between the image and the pseudo-tags obtained in (a), thus enhancing the image-text alignment.
@@
To address this, we shift from standard single image embedding to a new method focusing on the pixel nearest to each tag candidate. This is based on the finding that pixels closely correlated with a specific tag more accurately pinpoint image segments, as shown in Fig. 2(a). By utilizing the similarity between the nearest pixel and the tag as the measure for image-tag similarity, we mitigate bias and can accurately extract pseudo-tags that reflect the image-text relationship.
@@
Based on extracted pseudo-tags, we create valid foreground masks for all image-relevant tags, removing unnecessary tag masks. This led to a new self-distillation-based CLIP fine-tuning method, assuming optimal image-text relationships encompass all pseudo-tag contexts (Fig. 2(b)). We create a pseudo-label, a union of the similarity maps between the image and all pseudo-tags, making CLIP learn this as the real image-text map. This process allows the model to recognize all relevant tag regions, not just one, improving CLIP’s image-text alignment without additional annotations or models beyond CLIP.
@@
2 Related Work
CLIP [31] stands out as the most prominent model for image-text alignment. It possesses a remarkable ability to comprehend various free-form texts, making it widely utilized in zero-shot/open-vocabulary scenarios. Consequently, various CLIP variants [4, 41, 44, 25, 32] demonstrate comparable performance to models trained with task-specific annotations across many downstream tasks, such as classification and segmentation, even if they are trained with only image and text data.
Table 1: Comparison of Our Method with Related Approaches.
@@ 2.2.1 Utilization of External NLP models
Recent studies [23, 30, 38, 39, 24] have focused on extracting tags from text for learning, relying on external NLP models (Fig. 3(b)). While some studies utilize NLP parsers [26, 16, 1] to extract nouns, others leverage Large Language Models (LLM) like Vicuna [7]. However, these text-based tag selection techniques solely rely on text input and entirely disregard image information. This leads to two issues as shown in Fig. 5: 1) Extracting image-irrelevant tags. 2) Overlooking image-relevant tags, particularly non-nouns. In contrast, our method avoids these issues by exclusively utilizing both image and text for tag selection.
@@ 2.2.2 Additional Filtering using Image Information
Some models [37, 9] additionally consider the image to score tags. However, they utilize globally pooled image representation, suffering from single tag bias (Fig. 1(b)). CoCu [37] additionally introduces scoring based on retrieved image relevancy, but it depends on the images retrieved, where tag-relevant images lead to a low score and vice versa, and requires high computational cost (see Sec. 0.D.1). Also, both models rely on external NLP models and fail to address the omission of image-relevant tags.
@@
(a) Tag Selection by Pixel-Tag Scoring.
(b) Method Overview.
Figure 4: Overall Framework of Our Method. (a) Tag Selection by Pixel-Tag Scoring: Utilizing the similarity with the most correlated pixel as scoring, we identify a significant gap between tags related to the image and irrelevant ones. This gap guides us in selecting tags with the largest score gap as the threshold. (b) Method Overview: Our fine-tuning approach begins by obtaining pseudo-tags through our tag selection process, which are then used to create an ideal heatmap representing the image-text relationship. We address single tag bias by distilling information between the ideal and actual heatmaps. Additionally, we integrate an auxiliary loss to learn the pixel-tag relationship. Throughout this process, both the image and text encoders are trained.
@@ 3.1 Tag Selection by Pixel-Tag Scoring
Given an image $I$ of size $H\times W$ and corresponding text $T$ composed of tags $\mathcal{T}=\{t_{1}\cdots t_{n}\}$, our objective is to extract tags, $\mathcal{T}_{gt}\subset\mathcal{T}$ that accurately represent the image-text relationship. (e.g., for the first image in Fig. 5, $\mathcal{T}_{gt}=\{t_{1},t_{2},t_{3},t_{6},t_{7}\}$.) Importantly, our approach relies solely on the image $I$ and the text $T$, without any external datasets or models.
Let $\mathcal{E}_{V},\mathcal{E}_{T}$ denote the image and text encoders, respectively, which provide pixel embeddings and global text embeddings: $\mathcal{E}_{V}(I)\in\mathbb{R}^{H\times W\times C},\mathcal{E}_{T}(T)\in\mathbb{R}^{C}$ where $C$ is the embedding dimension size of the encoder output and $H,W$ are height and width of the image. The global image embedding can be obtained by pooling the pixel embedding: $\mathrm{pool}(\mathcal{E}_{V}(I))\in\mathbb{R}^{C}$.
@@
$s_{\mathrm{image}}(t_{i}):=\mathrm{sim}(\mathrm{pool}(\mathcal{E}_{V}(I)),\mathcal{E}_{T}(t_{i})),$ (1)
@@
Subsequently, we notice a substantial score gap between tags associated with the image and those not. As illustrated in Fig. 4(a), we arrange all tags according to $s_{\mathrm{pixel}}^{\mathrm{ours}}$ and compute the score difference between consecutive tags. We observed the significantly large score difference between meaningful tags and others at the boundary. Leveraging this observation, we extract pseudo-tags with high scores using the largest score gap as the boundary: $\mathcal{T}_{pred}\subset\mathcal{T}$.
@@ 3.2 Text-Tag Self-Distillation
In this section, we introduce a self-distillation technique designed to mitigate single tag bias arising from text embedding (Fig. 1(a)). The pseudo-tags $\mathcal{T}_{pred}$ obtained from our tag selection method (Sec. 3.1) encapsulate the relationship between the image and the text, enabling the transfer of knowledge from pseudo-tags to the text. Specifically, we ensure alignment between the union of similarity maps of pseudo-tags and the image with that of the text and image:
$\mathrm{simmap}(x):=\mathrm{sim}(\mathcal{E}_{V}(I),\mathcal{E}_{T}(x))\in\mathbb{R}^{H\times W},$ (3)
$\mathcal{L}_{distill}=||\mathrm{simmap}(T)-\bigcup_{t_{i}\in\mathcal{T}_{pred}}\mathrm{norm}(\mathrm{simmap}(t_{i}))||^{2}_{2},$ (4)
where $\mathrm{simmap}(x)$ is an abbreviation of the similarity map between the image $I$ and a text $x$, $\bigcup$ is the pixel-wise (element-wise) maximum operation, and $\mathrm{norm}$ is min-max normalization. Here, we apply a stop-gradient operation on the union similarity map to propagate the gradients only through the text-based one.
Additionally, we introduce an auxiliary loss function to reinforce the correlation between tags and images. Non-selected tags are deemed irrelevant to the image; thus, we drive their similarity map toward zero. Conversely, selected tags are trained to align with the discrete map to disregard unrelated image parts:
[…]
We also apply a stop-gradient operation on the normalized masks.
@@
Some may question whether it is more effective to view segment-level relationships using external models like SAM [19] rather than examining pixel-level relationships. However, our method is unrestricted and versatile since we self-train the model utilizing only image and text information without additional supervision, such as ground-truth tag annotation or external models. We also provide supplementary experiments on advanced techniques utilizing external models in Sec. 0.D.5.
@@ 4 Experiments
4.0.1 Datasets and Implementation Details
In all experiments, we train our model solely with image-text pairs without utilizing additional annotations like ground-truth tag information for the text. We demonstrate its flexibility by evaluating its performance on two distinct public models [44, 4]. Furthermore, to cut computational costs, we utilize the LoRA [17] technique for all our training processes. We use the CC3M and CC12M datasets [34, 5] for training, applying an additional sample filtering strategy to reduce the use of largely uncorrelated image-text pairs.
To evaluate the performance of our tag selection method, we manually collect ground-truth tag information for all image-text pairs of CC3M validation dataset [34]. Also, to see how much the single tag bias has eased, we generate the tag’s corresponding mask data for 680 image-text pairs. This is necessary because, to our knowledge, there is no existing dataset that provides a complete match of all ground-truth tags and their corresponding masks with the text. We will make this information available through code release. Details are in Appendix 0.E.
Additionally, for assessing our method’s effectiveness in open-vocabulary semantic segmentation using tag inputs, we use six established benchmarks: PASCAL VOC (VOC) [12], PASCAL Context (Context) [29], COCO-Stuff (Stuff), COCO-Object (Object) [3], Cityscapes (City) [10], and ADE20k (ADE) [43]. Further evaluations on segmentation performance with referring datasets, alongside conventional open-vocabulary classification tasks, and specifics of these assessments are detailed in Sec. 0.C.1 and 0.D.
@@ 4.1 Main Results
(a) Comparison based on the use of NLP models.
@@
(b) Comparison based on scoring methods.